A 25.00 mL sample of hydrochloric acid was titrated with 0.7425 M sodium hydroxide. A total of 21.24 mL of the base was required to reach the equivalence point.


a) Write a balanced molecular equation for this reaction?


b) What was the concentration of the hydrochloric acid?



1
Expert's answer
2021-05-27T07:34:52-0400

a)HCl + NaOH -->> NaCl + H2O

b) 0.7425/25 = 21.24/X

= 0.7425 ×21.24/25

= 0.631g/mol